A customer purchases a television for $800 using a credit card. The interest is charged on any unpaid balance at the rate of 18% per year compounded monthly. If the customer makes no payment for one year, how much is owed at the end of the year?
Answer by Theo(13342)

f = p * (1 + r) ^ n
f is the future value
p is the present value
r is he interest rate per time period
is the number of time periods.

in this problem, the time periods are in months.

the number of months is 1 year * 12 = 12 months.
the interest rate per month is 18% per year divided by 12 = 18/12% = 1.5% per month.

the formula becomes:

f = 800 * (1 + .015) ^ 12 = 956.4945372.

the amount of money owed at the end of 1 year is 956.49 rounded to the nearest penny.

in this formula, the rate is used, not the percent.
rate = percent / 100.